OpenStudy (anonymous):

Use elimination to solve the system of equations. Enter your answer as an ordered triple; example: (2, -3, 4) 4x+y+3z=0 2x-2y-z=10 3x-2y+2z=11

OpenStudy (anonymous):

8x+2y+6z=0 (from first equation) (1) 2x-2y-z=10 (2) 10x+5z=10 (1+2) ... (5) 2x-2y-z=10 (3) 3x-2y+2z=11 (4) x+3z=1 (4+3) ... (6) 10x+5z=10 (5) 10x+15z=10 (from (6)) ... (7) 10z=0 (7-5) z=0 10x+5(0)=10 10x=10 x=1 4x+y+3z=0 4(1)+y+3(0)=0 4+y=0 y=-4 (1,-4,0)